# Dynamics of Pseudoentanglement

### v1

@@ II.2 Equilibration of Markov chains and cut-off phenomenon
In many cases of interest in physics, these two time scales have the same behavior: e.g., for the simple random walk on a line of length $L$, both time scales are $\propto L^{2}$. This gives a scale-invariant behavior, with the approach to equilibrium depending on the single ratio $t/L^{2}$. However, in highly-connected spaces (e.g. expander graphs Hoory et al. (2006)) the two time scales can be different, with $t_{\rm rel}/t_{\rm mix}\to 0$ in the large-system limit. Such a separation of time scales, sketched in Fig. 2, is conjectured Peres (2004); Chen and Saloff-Coste (2008) to give rise to the cut-off phenomenon Diaconis and Shahshahani (1981); Diaconis (1996); Chen and Saloff-Coste (2008); Lacoin and Leblond (2011); Chen (2021): an abrupt transition to equilibrium, defined by
$\lim_{N\to\infty}\|p_{t(N)}-\pi\|_{\rm tv}=\chi(\theta),\quad t(N)=t_{\rm mix}(N)+\theta t_{\rm rel}(N).$ (9)
Here $\chi(\theta)$ is an $N$-independent function that goes to 1 as $\theta\to-\infty$ and to 0 as $\theta\to+\infty$. Informally, the total variation distance from equilibrium exhibits a scaling collapse as $N\to\infty$ on the form $\chi[(t-t_{\rm mix})/t_{\rm rel}]$, where the crucial observation is that the mid-point of the collapse ($t_{\rm mix}$) and its width ($t_{\rm rel}$) scale differently in the large-system limit: $t_{\rm rel}=o(t_{\rm mix})$. In other words, if we express time in units of $t_{\rm mix}$, the distance from equilibrium becomes step-like in the infinite-system limit, going from almost-maximal to almost-zero over a vanishingly narrow window of time, $1\pm t_{\rm rel}/t_{\rm mix}$. This is sketched in Fig. 2(b).
II.3 Random walks on permutation groups
In this work we will be especially interested in Markov chains over permutation groups and related processes. Here we give a brief review of some essential concepts Levin and Peres (2017).
@@ II.3 Random walks on permutation groups
The IP induces other Markov chains of interest. The random walk (RW) with initial state $v_{i}$ is the Markov chain $\{\sigma_{t}(v_{i})\}_{t=0}^{\infty}$. This has the intuitive meaning we expect for a random walk on the graph $(V,E)$: a particle initialized at vertex $v_{i}$ can hop with equal probability to any of its neighboring vertices. Additionally, the exclusion process with $m$ particles ($\textsf{EX}(m)$) is defined as the Markov chain $\{\sigma_{t}(A)\}_{t=0}^{\infty}$ where $A\subset V$ is a collection of $m$ distinct vertices, and $\sigma_{t}(A)\equiv\{\sigma_{t}(v_{i}):v_{i}\in A\}$ is its image under permutation $\sigma_{t}$. This process describes the evolution of $m$ indistinguishable, hard-core particles on the graph, where at each time step the occupation number of two neighboring vertices get swapped. When $m=1$, this clearly reduces to the random walk: $\textsf{EX}(m)=\textsf{RW}$.
These three Markov chains are connected by a remarkable result known as Aldous’ spectral gap conjecture [^3] Despite the name, this is a proven theoremCaputo et al. (2010)., which states that
@@
The inequality $t_{\rm rel}^{\textsf{IP}}\geq t_{\rm rel}^{\textsf{EX}(m)}\geq t_{\rm rel}^{\textsf{RW}}$ is a simple consequence of the fact that RW and $\textsf{EX}(m)$ are sub-processes of IP Caputo et al. (2010) (informally, they are projections of IP onto a smaller state space), and thus it holds also for other Markov chains on $\mathcal{S}_{|V|}$ beyond IP. However, the other direction of the inequality is highly nontrivial and is specific to IP, where it was conjectured in 1992 and proved in 2009 Caputo et al. (2010).
Beyond IP, in this work we will consider other Markov chains over permutation groups. In particular, we will consider a setting where the unit time step does not just transpose two neighboring vertices, like IP, but rather performs one of a more general set of “elementary permutations” in $\mathcal{S}_{|V|}$ sampled randomly at each time step. Such a process also gives a Markov chain over $\mathcal{S}_{|V|}$ and also induces “generalized exclusion processes” through the action of $\mathcal{S}_{|V|}$ on the subset spaces $\Sigma_{m}$, in the same way as IP induces $\textsf{EX}(m)$. These “generalized exclusion processes” differ from the standard one as they can allow multiple particles to hop in a correlated way in each time step. Aldous’ spectral gap conjecture, Eq. (10), does not apply to this more general case. In Sec. IV.3 we will in fact encounter a case where the $m$-particle “generalized exclusion process” relaxes more slowly than the single-particle walk.
III Pseudothermalization in automaton circuits
Figure 3: (a) Subset states and their transformation under automaton gates, illustrated here for $N=3$ qubits and a CCX gate. (b) Spaces of subsets of the computational basis, $\Sigma_{m}=\{S\subset\mathbb{Z}_{2}^{N}:\,|S|=m\}$, shown for $N=3$ and $m=1,2,3$. Vertices represent subsets $S\in\Sigma_{m}$, edges represent the action of an automaton gate set $\mathcal{G}$ (self-edges are omitted for simplicity).
We aim to introduce models of quantum dynamics that achieve pseudoentanglement as their equilibrium state. However, fully-generic quantum dynamics tends to maximize actual entanglement. The easiest way to avoid this fate is to constrain the dynamics to preserve the ‘subset(-phase) state’ property, Sec. II.1. Reversible classical circuits, also known as automaton circuits in the quantum context [^4] Note that some works use a different definition of automaton gates, which includes phases Iaconis (2021). Gopalakrishnan (2018); Iaconis et al. (2019); Iaconis (2021); Klobas et al. (2021), have this property: an automaton circuit $U_{t}$ corresponds to a permutation $\sigma_{t}\in\mathcal{S}_{D}$ of the computational basis, $U_{t}\ket{\mathbf{z}}\equiv\ket{\sigma_{t}\mathbf{z}}$; thus, its action on a subset-phase state |$\ket{\psi_{S_{0},f_{0}}}$ is
@@ III Pseudothermalization in automaton circuits
In this work we consider the following type of dynamics. A distribution of initial states |$\ket{\psi_{S_{0},f_{0}}}$ is specified (these may be subset states or subset-phase states depending on the allowed values of $f_{0}$); then, each step of the time evolution consists of sampling an automaton gate $u_{i}$ uniformly at random from a fixed gate set $\mathcal{G}$ (we will discuss requirements on $\mathcal{G}$ in the following) and applying it to the state. After $t$ time steps, this results in an ensemble of time-evolved states $\{U_{t}\ket{\psi_{S_{0},f_{0}}}=\ket{\psi_{S_{t},f_{t}}}\}$ defined by all possible initial states and choices of $U_{t}=u_{i_{t}}u_{i_{t-1}}\cdots u_{i_{1}}\in\mathcal{G}^{\circ t}$. The relevant notion of equilibration for this state ensemble is whether, at late times, it becomes uniformly distributed on the appropriate (subset or subset-phase) state space. If that is the case, and subject to usual constraints on the subset size $K$ (see Sec. II.1), then our dynamics achieves pseudoentanglement at late times. We call this condition pseudothermalization: while the upper limit $K$ on the Schmidt rank prevents full thermalization (defined, for a model without conservation laws such as this one, by the approach to a Page state), this failure of thermalization may be practically undetectable, in a sense that is rigorously formalized through the concept of pseudoentanglement.
A first question to address is whether such equilibrium is achieved at all for a given automaton gate set $\mathcal{G}$. Let us focus on the subsets and neglect the phases for now. The action of the quantum circuits $U_{t}\ket{\psi_{S_{0}}}=\ket{\psi_{S_{t}}}$ defines a Markov chain $\{S_{t}\}$ over the space of size-$K$ subsets of the computational basis,
@@
$\Gamma_{S,S^{\prime}}=\frac{1}{|\mathcal{G}|}\sum_{u\in\mathcal{G}}\delta_{S,u(S^{\prime})}.$ (13)
Viewing $\Gamma$ as an adjacency matrix, we can thus picture the $\Sigma_{K}$ spaces as graphs; some minimal examples are shown in Fig. 3(b). To achieve pseudothermalization, the Markov chain $\{S_{t}\}$ must converge at late times to the uniform distribution on $\Sigma_{K}$, denoted here as $\pi_{K}$:
@@
where $p_{t}(S)$ is the probability distribution over subsets after $t$ time steps. To guarantee the uniform distribution $\pi_{K}$ as a unique steady state, the Markov chain must be Levin and Peres (2017):
@@
    Reversible: the Markov chain transition matrix is symmetric, $\Gamma_{S^{\prime},S}=\Gamma_{S,S^{\prime}}$ for all $S,S^{\prime}$. This property is ensured e.g. by picking a gate set $\mathcal{G}$ such that $u^{2}=I$ for all $u\in\mathcal{G}$.
@@
    Irreducible: for any two states $S,S^{\prime}\in\Sigma_{K}$, there exists a sequence of gates in $\mathcal{G}$ that maps $S$ to $S^{\prime}$. Irreducibility will be discussed separately for each gate set $\mathcal{G}$ in the following.
@@
    Aperiodic: the chain can connect any two states $S,S^{\prime}$ with paths of any (sufficiently large) length [^5] The prototypical counterexample is given by walks on bipartite graphs, where all paths connecting $S$ to $S^{\prime}$ have the same length modulo 2. In that case the populations of the two partitions can keep oscillating indefinitely.. One way to ensure aperiodicity is to make the chain ‘idle’, i.e. allow it to skip a step. This can be done by including the identity gate in $\mathcal{G}$.
For a gate set $\mathcal{G}$ with all three properties above, we can conclude that $p_{t}\to\pi_{K}$ as $t\to\infty$. This suffices to establish pseudoentanglement at late times in two cases:
  • when the initial state distribution is restricted to a constant phase function, e.g. $f_{0}=0$, we obtain the random subset ensemble Giurgica-Tiron and Bouland (2023);
  • when the initial state distribution includes a uniformly random phase function $f_{0}$, we obtain the random subset-phase ensemble Aaronson et al. (2023).
Both of these ensembles are pseudoentangled, subject to certain constraints on $K$, as we reviewed in Sec. II.1.
Having established that the steady-state ensembles in these two cases are pseudoentangled, it remains to understand how the system approaches this condition over the course of its dynamics, and what are the time scales involved. As already mentioned, we will call this process pseudothermalization, in analogy with how conventional thermalization in quantum many-body systems can be defined from the saturation of entanglement of local subsystems. Two key technical results will allow us to make progress on these questions by connecting pseudothermalization with the equilibration of certain classical Markov chains.
@@
$\left\|\rho^{(m)}_{\mathcal{E}_{p}}-\rho^{(m)}_{\rm Haar}\right\|_{\rm tr}=2\|\Phi_{m\leftarrow K}[p]-\pi_{m}\|_{\rm tv}+O\left(\frac{m^{2}}{K}\right)$ (15)
Here $\|\cdots\|_{\rm tr}$ and $\|\cdots\|_{\rm tv}$ are the trace norm and total variation norm, respectively; $\rho^{(m)}$ denotes the $m$-th statistical moment of a state ensemble, i.e. $\rho^{(m)}_{\mathcal{E}_{p}}=\mathbb{E}_{S\sim p}\mathbb{E}_{f}[\outerproduct{\psi_{S,f}}{\psi_{S,f}}^{\otimes m}]$ and $\rho^{(m)}_{\rm Haar}=\int{\rm d}\psi_{\rm Haar}\outerproduct{\psi}{\psi}^{\otimes m}$; and $\Phi_{m\leftarrow K}$ is a stochastic map from distributions on $\Sigma_{K}$ to distributions on $\Sigma_{m}$, defined by
@@
If we take $m={\rm poly}(N)$ and $K=\omega({\rm poly}(N))$, then the error term on the RHS of Eq. (15) is negligible, and pseudoentanglement (defined by the LHS being small) is tied to the non-uniformity of the distribution $\Phi_{m\leftarrow K}[p]$ on $\Sigma_{m}$. If we view $p$ as the distribution $\Sigma_{K}$ induced by the action of a random automaton circuit, then $\Phi_{m\leftarrow K}[p]$ is the distribution induced by the same process on $\Sigma_{m}$ (up to an average over initial states). We show this in Appendix B, along with other useful properties of the stochastic maps $\Phi$. Then Eq. (15) maps pseudoentanglement onto the mixing of a classical Markov chain on $\Sigma_{m}$.
@@
$\left\|\rho_{\mathcal{E}^{\prime}_{p}}^{(m)}-\rho_{\rm Haar}^{(m)}\right\|_{\rm tr}=\left\|\mathcal{M}^{(m)}_{p}\right\|_{\rm tr}+O\left(\frac{m}{\sqrt{K}}\right).$ (17)
@@
where we use subsets $S^{\prime},S^{\prime\prime}\in\Sigma_{m}$ to index the matrix elements, and define $\delta\equiv|S^{\prime}\setminus S^{\prime\prime}|$.
@@
We again relate pseudoentanglement (the LHS of Eq. (17) being small) to the non-uniformity of certain probability distributions over subsets; except unlike the previous case, Eq. (15), here we have contributions from a family of distributions $\{\Phi_{m+\delta\leftarrow K}[p]\}_{\delta=0}^{m}$, rather than from the single distribution $\Phi_{m\leftarrow K}[p]$. If we view $p$ as arising from a Markov chain on $\Sigma_{K}$ induced by a random automaton circuit dynamics, then the emergence of pseudoentanglement is again related—albeit in a less straightforward way—to the mixing of the associated Markov chains $\Phi_{m+\delta\leftarrow K}[p]$ on $\Sigma_{m+\delta}$, for all $\delta=0,1,\dots,m$.
@@
Having reduced the problem of pseudothermalization to that of equilibration of certain Markov chains (namely random walks over the subset spaces $\{\Sigma_{m}\}$, Fig. 3(b)), we are now in a position to study the problem quantitatively by leveraging known results on these processes, including those reviewed in Sec. II.2 and II.3. In the following sections we consider several automaton gate sets $\mathcal{G}$ and initial state distributions, and analyze the associated classical Markov chains to characterize the process of pseudothermalization.
IV Spreading of Pseudoentanglement
IV.1 Motivation and setup

### v3

@@ 2.2 Equilibration of Markov chains and cut-off phenomenon
In many cases of interest in physics, these two time scales have the same behavior: e.g., for the simple random walk on a line of length $L$, both time scales are $\propto L^{2}$. This gives a scale-invariant behavior, with the approach to equilibrium depending on the single ratio $t/L^{2}$. However, in some cases (including random walks on expander graphs hoory_expander_2006 , which are highly connected spaces) the two time scales can be different, with $t_{\rm rel}/t_{\rm mix}\to 0$ in the large-system limit. Such a separation of time scales, sketched in Fig. 2, is conjectured peres_conjecture ; chen_cutoff_2008 to give rise to the cut-off phenomenon diaconis_generating_1981 ; diaconis_cutoff_1996 ; chen_cutoff_2008 ; lacoin_cutoff_2011 ; chen_cutoff_2021 : an abrupt transition to equilibrium, defined as
$\lim_{N\to\infty}D_{\rm tv}(p_{t_{\rm mix}(N)+\theta t_{\rm rel}(N)},\pi)$ $=\chi(\theta).$ (10)
Here $\chi(\theta)$ is an $N$-independent function that goes to 1 as $\theta\to-\infty$ and to 0 as $\theta\to+\infty$. Informally, as $N\to\infty$, the total variation distance from equilibrium exhibits a scaling collapse on the form $\chi[(t-t_{\rm mix})/t_{\rm rel}]$, where crucially the mid-point of the collapse ($t_{\rm mix}$) and its width ($t_{\rm rel}$) scale differently in the large-system limit, $t_{\rm rel}=o(t_{\rm mix})$. Thus if we express time in units of $t_{\rm mix}$, the distance from equilibrium becomes step-like in the infinite-system limit, going from almost-maximal to almost-zero over a vanishingly narrow window of time, $1\pm t_{\rm rel}/t_{\rm mix}$. This is sketched in Fig. 2(b).
2.3 Random walks on permutation groups
In this work we will be especially interested in Markov chains over permutation groups and related processes. Here we give a brief review of some essential concepts; for a more thorough discussion, see e.g. Ref. levin_markov_2017 .
@@ 2.3 Random walks on permutation groups
The IP induces other Markov chains of interest. The random walk (RW) with initial state $v_{i}$ is the Markov chain $\{\sigma_{t}(v_{i})\}_{t=0}^{\infty}$. This has the intuitive meaning we expect for a random walk on the graph $(V,E)$: a particle initialized at vertex $v_{i}$ can hop with equal probability to any of its neighboring vertices (or stay in place). Additionally, the exclusion process with $m$ particles ($\textsf{EX}(m)$) is defined as the Markov chain $\{\sigma_{t}(A)\}_{t=0}^{\infty}$ where $A\subset V$ is a collection of $m$ distinct vertices, and $\sigma_{t}(A)\equiv\{\sigma_{t}(v_{i}):v_{i}\in A\}$ is its image under permutation $\sigma_{t}$. This process describes the evolution of $m$ indistinguishable, hard-core particles on the graph, where at each time step the occupation number of two neighboring vertices get swapped. When $m=1$, this clearly reduces to the random walk: $\textsf{EX}(1)=\textsf{RW}$.
These three Markov chains (RW, $\textsf{EX}(m)$ and IP) are connected by a remarkable result known as Aldous’ spectral gap conjecture [^8] Despite the name, this is a proven theoremcaputo_proof_2010 ., which states that
@@
The inequalities $t_{\rm rel}^{\textsf{IP}}\geq t_{\rm rel}^{\textsf{EX}(m)}\geq t_{\rm rel}^{\textsf{RW}}$ are a simple consequence of the fact that $\textsf{EX}(m)$ is a sub-process of IP caputo_proof_2010 (informally, a projection of IP onto a smaller state space), and that RW is a subprocess of $\textsf{EX}(m)$ ($m>1$). However, the other direction of the inequality, $t_{\rm rel}^{\textsf{IP}}\leq t_{\rm rel}^{\textsf{EX}(m)}\leq t_{\rm rel}^{\textsf{RW}}$, is specific to IP and is highly nontrivial—it was conjectured in 1992 and proved in 2009 caputo_proof_2010 .
Beyond IP, in this work we will consider other Markov chains over permutation groups. In particular, we will consider a setting where the unit time step does not just transpose two neighboring vertices, like in IP, but rather performs one of a more general set of “elementary permutations” in $\mathcal{S}_{|V|}$ sampled randomly at each time step. Such a process also gives a Markov chain over $\mathcal{S}_{|V|}$ and also induces “generalized exclusion processes” through the action of $\mathcal{S}_{|V|}$ on the subset spaces $\Sigma_{m}$, in the same way as IP induces $\textsf{EX}(m)$. These “generalized exclusion processes” differ from the standard one as they can allow multiple particles to hop in a correlated way in each time step. Aldous’ spectral gap conjecture, Eq. (11), does not apply to these more general cases. In Sec. 4.3 we will in fact encounter a case where the $m$-particle “generalized exclusion process” relaxes more slowly than the single-particle random walk.
3 Ensemble pseudothermalization in automaton circuits
Figure 3: (a) Subset states and their transformation under automaton gates, illustrated here for $N=3$ qubits and a CCX gate. (b) Spaces of subsets of the computational basis, $\Sigma_{m}=\{S\subset\mathbb{Z}_{2}^{N}:\,|S|=m\}$, shown for $N=3$ and $m=1,2,3$. Vertices represent subsets $S\in\Sigma_{m}$, edges connect subsets $S,S^{\prime}$ that are mapped to each other under an element of the automaton gate set $\mathcal{G}$ (self-edges are omitted for simplicity).
3.1 Criteria for ensemble pseudothermalization
We aim to introduce models of quantum dynamics that achieve pseudoentanglement as their equilibrium state. However, fully-generic quantum dynamics tends to maximize actual entanglement. The easiest way to avoid this fate is to constrain the dynamics to preserve the ‘subset(-phase) state’ property, Sec. 2.1. Reversible classical circuits, also known as automaton circuits in the quantum context [^9] Note that some works use a slightly different definition of automaton gates which also includes phases iaconis_quantum_2021 . gopalakrishnan_operator_2018 ; iaconis_anomalous_2019 ; iaconis_quantum_2021 ; klobas_exact_2021 , have this property: an automaton circuit $U_{t}$ corresponds to a permutation $\sigma_{t}\in\mathcal{S}_{D}$ of the computational basis, $U_{t}\ket{\mathbf{z}}\equiv\ket{\sigma_{t}(\mathbf{z})}$; thus, its action on a subset-phase state |$\ket{\psi_{S_{0},f_{0}}}$ is
@@ 3.1 Criteria for ensemble pseudothermalization
In this work we consider the following type of dynamics. A distribution of initial states |$\ket{\psi_{S_{0},f_{0}}}$ is specified (these may be subset states or subset-phase states depending on the allowed values of $f_{0}$); then, each step of the time evolution consists of sampling an automaton gate $u_{i}$ uniformly at random from a fixed gate set $\mathcal{G}$ and applying it to the state. (We will discuss requirements on the gate set $\mathcal{G}$ in the following.) After $t$ time steps, this results in an ensemble of time-evolved states $\{U_{t}\ket{\psi_{S_{0},f_{0}}}=\ket{\psi_{S_{t},f_{t}}}\}$ defined by all possible initial states {|$\{\ket{\psi_{S_{0},f_{0}}}\}$ and all possible choices of $U_{t}=u_{i_{t}}u_{i_{t-1}}\cdots u_{i_{1}}\in\mathcal{G}^{\circ t}$. The relevant notion of equilibration for this state ensemble is whether, at late times, it becomes uniformly distributed on the appropriate (subset or subset-phase) state space. If that is the case, and subject to usual constraints on the subset size $K$ (see Sec. 2.1), then our dynamics achieves pseudoentanglement at late times. This is the condition we termed ensemble pseudothermalization in Sec. 1: while the upper limit $K$ on the Schmidt rank prevents maximization of the entanglement entropy and equilibration to a Page state, this departure from the predictions of quantum thermalization may be practically undetectable, in a sense that is rigorously formalized through the concept of pseudoentanglement.
A first question to address is whether such equilibrium is achieved at all for a given automaton gate set $\mathcal{G}$. Let us focus on the subsets and neglect the phases for now. The action of automaton circuits on subset states, $U_{t}\ket{\psi_{S_{0}}}=\ket{\psi_{S_{t}}}$, defines a Markov chain $\{S_{t}\}$ over the space of size-$K$ subsets of the computational basis,
@@
$\Gamma_{S^{\prime},S}=\frac{1}{|\mathcal{G}|}\sum_{u\in\mathcal{G}}\delta_{S^{\prime},u(S)}.$ (14)
This transition matrix represents the fact that each gate $u\in\mathcal{G}$ is chosen with uniform probability $1/|\mathcal{G}|$, and maps each subset $S$ to a subset $S^{\prime}=u(S)\equiv\{u(\mathbf{z}):\,\mathbf{z}\in S\}$ (possibly the same subset). Viewing $\Gamma$ as an adjacency matrix, we can thus picture the $\Sigma_{K}$ spaces as graphs. Some minimal examples are shown in Fig. 3(b), and reveal substantial complexity already for very small systems and number of copies considered. To achieve pseudoentanglement, the Markov chain $\{S_{t}\}$ must converge at late times to the uniform distribution on $\Sigma_{K}$, denoted here as $\pi_{K}$:
@@
where $p_{t}(S)$ is the probability distribution over subsets after $t$ time steps.
To guarantee the uniform distribution $\pi_{K}$ as a unique steady state, the Markov chain must be levin_markov_2017 :
@@
    Reversible—the Markov chain transition matrix is symmetric, $\Gamma_{S^{\prime},S}=\Gamma_{S,S^{\prime}}$ for all $S,S^{\prime}$. This property is ensured, e.g., by picking a gate set $\mathcal{G}$ such that $u^{2}=I$ for all $u\in\mathcal{G}$.
@@
    Irreducible—for any two states $S,S^{\prime}\in\Sigma_{K}$, there exists a sequence of gates in $\mathcal{G}$ that maps $S$ to $S^{\prime}$. Irreducibility will be discussed separately for each gate set $\mathcal{G}$ in the following.
@@
    Aperiodic—the chain can connect any two states $S,S^{\prime}$ with paths of any (sufficiently large) length [^10] The prototypical counterexample is given by non-idle walks on bipartite graphs, where all paths connecting $S$ to $S^{\prime}$ have the same length modulo 2. In that case the populations of the two partitions of the graph can keep oscillating indefinitely preventing convergence to a steady state.. One way to ensure aperiodicity is to make the chain ‘idle’, i.e., allow it to skip a step with some probability. This can be done by including the identity gate in $\mathcal{G}$.
For a gate set $\mathcal{G}$ with all three properties above, we can conclude that $p_{t}\to\pi_{K}$ as $t\to\infty$ levin_markov_2017 . This suffices to establish pseudoentanglement at late times in two cases:
  • when the initial state distribution is restricted to a constant phase function, e.g. $f_{0}=0$, we obtain the random subset ensemble giurgica-tiron_pseudorandomness_2023 ; jeronimo_pseudorandom_2024 ;
  • when the initial state distribution includes a uniformly random phase function $f_{0}$, we obtain the random subset-phase ensemble aaronson_quantum_2023 .
Both of these ensembles are pseudoentangled, subject to certain constraints on $K$, as we reviewed in Sec. 2.1.
3.2 Mapping to classical Markov processes
Having established that the steady-state ensembles in these two cases are pseudoentangled, it remains to understand how the system approaches this condition over the course of its dynamics, and what are the time scales involved. As already mentioned, we will call this process ensemble pseudothermalization, in analogy with how conventional thermalization in quantum many-body systems can be defined from the saturation of entanglement of local subsystems [^11] Pseudoentanglement however requires an ensemble of states, hence ‘ensemble’ pseudothermalization. Two key technical results will allow us to make progress on these questions by connecting pseudoentanglement with the equilibration of certain classical Markov chains.
@@ 3.2 Mapping to classical Markov processes
$D_{\rm tr}\left(\rho^{(m)}_{\mathcal{E}_{p}},\rho^{(m)}_{\rm Haar}\right)=D_{\rm tv}\left(\Phi_{m\leftarrow K}[p],\pi_{m}\right)+O\left(\frac{m^{2}}{K}\right)$ (16)
Here $D_{\rm tr}$ and $D_{\rm tv}$ are the trace distance between quantum states and total variation distance between probability distributions, respectively; $\rho^{(m)}$ denotes the average $m$-copy state of an ensemble, i.e. $\rho^{(m)}_{\mathcal{E}_{p}}=\mathbb{E}_{S\sim p}\mathbb{E}_{f}[\outerproduct{\psi_{S,f}}{\psi_{S,f}}^{\otimes m}]$ and $\rho^{(m)}_{\rm Haar}=\int{\rm d}\psi_{\rm Haar}\outerproduct{\psi}{\psi}^{\otimes m}$; and $\Phi_{m\leftarrow K}$ is a stochastic map from distributions on $\Sigma_{K}$ to distributions on $\Sigma_{m}$, defined by
@@
This statement quantitatively links the distinguishability between quantum ensembles [left hand side of Eq. (16)] to the (non-)uniformity of a probability distribution over $\Sigma_{m}$, that is, over subsets of the computational basis of cardinality $m$ [right hand side of Eq. (16)]. Since we take $m\leq O({\rm poly}(N))$ and $K\geq\omega({\rm poly}(N))$, the error term on the right hand side is negligible in the regime we are interested in. Note that the ensemble $\mathcal{E}_{p}$ is determined by a probability distribution $p$ over $\Sigma_{K}$, but what controls the $m$-copy distinguishability is a different distribution, $\Phi_{m\leftarrow K}[p]$, over $\Sigma_{m}$. We may view $\Sigma_{K}$ as the state space for $K$ indistinguishable particles on the hypercube $\{0,1\}^{N}$; then the stochastic map $\Phi_{m\leftarrow K}$ is akin to taking the marginal distribution over $m$-particle states. Particle indistinguishability however requires symmetrizing over all $\binom{K}{m}$ $m$-particle subsets of each $K$-particle subset, which is the content of Eq. (17).
In Appendix B we prove various useful properties of the stochastic maps $\Phi$. Among these is the fact that, if we view $p$ as the distribution on $\Sigma_{K}$ induced by the action of a random automaton circuit, then $\Phi_{m\leftarrow K}[p]$ is the distribution induced by the same process on $\Sigma_{m}$, up to an average over initial states. Then Eq. (16) maps the $m$-copy distinguishability between quantum state ensembles onto the mixing of a classical Markov chain on $\Sigma_{m}$.
@@
$D_{\rm tr}\left(\rho_{\mathcal{E}^{\prime}_{p}}^{(m)},\rho_{\rm Haar}^{(m)}\right)=\frac{1}{2}\left\|\mathcal{M}^{(m)}_{p}\right\|_{\rm tr}+O\left(\frac{m}{\sqrt{K}}\right).$ (18)
@@
where we use subsets $S^{\prime},S^{\prime\prime}\in\Sigma_{m}$ to index the matrix elements of $\mathcal{M}^{(m)}_{p}$, and define $\delta\equiv|S^{\prime}\setminus S^{\prime\prime}|$.
@@
This statement again relates distinguishability between quantum state ensembles [left hand side of Eq. (18)] to the non-uniformity of certain probability distributions over subsets [right hand side of Eq. (18)]. However, unlike Eq. (16), here we have contributions from a family of distributions $\{\Phi_{m+\delta\leftarrow K}[p]\}_{\delta=0}^{m}$, rather than from the single distribution $\Phi_{m\leftarrow K}[p]$. If we view $p$ as arising from a Markov chain on $\Sigma_{K}$ induced by a random automaton circuit dynamics, then the emergence of pseudoentanglement is again related—albeit in a less straightforward way—to the mixing of the associated Markov chains $\Phi_{m+\delta\leftarrow K}[p]$ on $\Sigma_{m+\delta}$, for all $\delta=0,1,\dots,m$.
@@
Having reduced the problem of ensemble pseudothermalization to that of equilibration of certain Markov chains [random walks over the subset spaces $\{\Sigma_{m}\}$, Fig. 3(b)], we are now in a position to study the problem quantitatively by leveraging known results on these processes, including those reviewed in Sec. 2.2 and 2.3. In the following sections we consider several automaton gate sets $\mathcal{G}$ and initial state distributions, and analyze the associated classical Markov chains to characterize the process of ensemble pseudothermalization.
4 Spreading of Pseudoentanglement
4.1 Motivation and setup